# A New Fick-Jacobs Derivation with Applications to Computational Branched Diffusion Networks

### v1

@@ 1 Introduction
The structure of the paper is as follows: in Section 2, we present the original derivation of the Fick-Jacobs model and introduce a modified derivation method that leads to our new model. In Section 3, we describe the one-dimensional spatial discretization methods we use for the different terms of our model, and how those discretizations are specialized for branch nodes in a one-dimensional network. In Section 4, we provide a stability analysis of our discretizations and prove new stability conditions on branching one-dimensional networks. Finally, in Section 5, we demonstrate that our method outperforms existing methods with respect to computational error, stability, and efficiency on multiple domains of changing radius, including a branched domain.
@@ 5.4 Branched Domain with Lateral Flux
We lastly consider a branched domain with lateral fluxes. This is something the classical Fick-Jacobs model was not designed to do, evident from eq. (1) having no term for boundary fluxes. In Jacobs’s original derivation, a key assumption was no boundary fluxes. Here, we use our expanded flux model to incorporate these.
@@ 6 Conclusion
Put together, we have established a computationally efficient and accurate dimension reduced model of diffusion through a network of porous channels. This has a wide array of applications in biology, engineering, and various other sciences. Our further interest lies specifically in neuroscience, where we aim to develop extended simulations of ion transport within neurons [10, 8, 7, 24, 26, 28, 6]. The dimension reduced model we have developed will help create long time simulations that give insight into the clinical regime of timescales, further assisting the study the effects of neurological treatments.

### v2

@@ 1 Introduction
The structure of the paper is as follows: in Section 2, we present the original derivation of the Fick-Jacobs model and introduce a modified derivation method that leads to our new model, termed Extended-Flux Fick-Jacobs (EF-FJ) model. In Section 3, we describe the one-dimensional spatial discretization methods we use for the different terms of our model, and how those discretizations are specialized for branch nodes in a one-dimensional network. In Section 4, we provide a stability analysis of our discretizations and prove new stability conditions on branching one-dimensional networks. Finally, in Section 5, we demonstrate that our method outperforms existing methods with respect to computational error, stability, and efficiency on multiple domains of changing radius, including a branched domain. Given the importance of the developed EF-FJ model to neuroscience we end with an example using a reconstructed neuron and simulate intracellular calcium dynamics inside the neuron.
@@ 5.4 Branched Domain with Lateral Flux
We next consider a branched domain with lateral fluxes. This is something the classical Fick-Jacobs model was not designed to do, evident from eq. (1) having no term for boundary fluxes. In Jacobs’s original derivation, a key assumption was no boundary fluxes. Here, we use our expanded flux model to incorporate these.
@@
5.5 An application to neuroscience
Lastly, we consider a neuron and the calcium dynamics in response to synaptic stimuli, which is mathematically modeled as diffusion through a tubular network. Calcium dynamics in neurons are important for the study of treatments for various neurological conditions, as calcium acts as a messenger in neurons, regulating processes such as neurotransmitter release, synaptic plasticity, neuronal excitability, and regulation of DNA transcription. Biologists and neuroscientists rely on accurate simulations to conduct reliable analysis and develop effective treatments. For this reason, we provide this as an essential application of our model.
We compare the results of our model to those of a simplified version of the model for calcium dynamics in [6],
$\frac{\partial c}{\partial t}$ $=$ $\frac{\partial}{\partial x}\left(D\frac{\partial c}{\partial x}\right)+\frac{2}{R}J_{PM},$ (27)
$J_{PM}$ $=$ $-J_{P}-J_{N}+J_{SYN}+J_{l,p}.$ (28)
For the reduced model (herein referred to as the Borole et al. model), we have removed consideration of the endoplasmic reticulum and buffering ionic reactions. Note that this model is equivalent to our expanded flux model as $\Delta x$, $dR/dx\to 0$. In the plasma membrane flux term, $J_{PM}$, we have included: plasma membrane Ca2+-ATPase (PMCA) pumps, $J_{P}$; Na+/Ca2+ exchangers (NCX), $J_{N}$; calcium influx through post-synaptic density, $J_{SYN}$; and the leak calcium flux across the plasma membrane, $J_{l,p}$. Details of the complete mathematical models for each of the $J_{PM}$ flux terms can be found in [6].
The complete neuronal geometry is shown in Fig. 14a. We select three nodes at which to provide a synaptic input: the soma (node S) and two branching nodes (nodes A and B). At each of these nodes, a 2 millisecond (ms) stimulus is applied every 10 ms, with the stimuli staggered by 0.5 ms (node S receives its first stimulus at 0 ms, node A at 0.5 ms, and node B at 1 ms). This periodic stimulus serves as a model for the effects of repetitive transcranial magnetic stimulation (rTMS) treatments on intracellular calcium dynamics. Additionally, we record the concentration at nearby nodes S1, A1, and B1, shown in Fig. 14b-14d, to compare the results of our model with the standard one in regions of large radial gradients.
Figure 14: Nodes (S, A, B) and nearby nodes (S1, A1, B1). (a) The complete neuronal model used with the radius indicated, and regions of each measuring node highlighted. (b)-(d) Zoom in of each highlighted region, detailing the region local to each measuring node and labels of node locations.
In Fig. 14a, we show the radius of the dendrites of the neuron in the region of consideration. We see a large radius at the soma (near node S), and a quick decrease in radius at each branching point, particularly clear near node B. ($dR/dx|_{S}=0.43$, $dR/dx|_{A}=-0.03$, $dR/dx|_{B}=-0.10$.) As will be demonstrated in this section, our expanded model computes significantly different concentrations in these regions in response to the changing radius.
When using a realistic geometry, satisfying stability conditions for the first order term is a significant obstacle. Central differencing is instead used for the first order term, which eases the requirements for stability. The same physical conservation considerations are taken at branching nodes as used in the derivation of the upwinding method. For additional stability in time stepping, a second order semi-implicit backward differentiation formula (SBDF2) is used [25].
In Fig. 15, we show the calcium concentration at each selected node using both the Borole et al. model and our EF-FJ model. The most prominent difference can be seen at node S, where the soma has a very significant radial change from the dendrite. There is a very slight difference in nodes A and B, where there is a large change in the radius at the branching node but very little change through the rest of the dendrite. We see particularly at nearby nodes S1, A1, and B1, a significant difference in results after the steep change in radius through the soma or past the branching points.
The most drastic difference is clearly seen at Node S, where, even though the stimulus is applied in the same way for both models, the drastic change in radius gives much different quantitative behavior. Our EF-FJ model shows a concentration approaching a relatively stable oscillation quickly, while the Borole et al. model continues to increase. This further demonstrates that the radial gradient consideration in our model has a significant impact on results and can provide more accurate results without the need for expensive higher dimensional computation.
(a) Node S
(b) Node S1
(c) Node A
(d) Node A1
(e) Node B
(f) Node B1
Figure 15: Concentration at each node (S, A, B) and nearby nodes (S1, A1, B1) using the EF-FJ model vs. Borole et al. model.
@@ 6 Conclusion
Put together, we have established a computationally efficient and accurate dimension reduced model of diffusion through a network of porous channels. This has a wide array of applications in biology, engineering, and various other sciences. Our further interest lies specifically in neuroscience, where we aim to develop extended simulations of ion transport within neurons [10, 8, 7, 24, 27, 29, 6]. The dimension reduced model we have developed will help create long time simulations that give insight into the timescales relevant for clinical applications.